An acne specialist has found that 60% of acne sufferers are helped by existing treatments. He decides to test a new treatment on a random sample of 40 acne patients. If 28 of these respond to this treatment, can he conclude that it is significantly better at the 0.05 level?


1
Expert's answer
2021-06-10T05:23:18-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0: p\leq0.6"

"H_1:p>0.6"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a right-tailed test is "z_c=1.6449."

The rejection region for this right-tailed test is "R=\{z: z>1.6449\}"


The z-statistic is computed as follows:


"\hat{p}=\dfrac{x}{n}=\dfrac{28}{40}=0.7"


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.7-0.6}{\sqrt{\dfrac{0.6(1-0.6)}{40}}}\approx1.291"

Since it is observed that "z=1.291<1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is greater than 0.6, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=P(z>1.291)=0.09835," and since "p=0.09835>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is greater than 0.6, at the "\alpha=0.05" significance level.


Therefore he cannot conclude that it is significantly better at the 0.05 level.